This is a goal-less problem. You can analyze both circuits or just one, your goal is just to demonstrate that your ability toWhich way the electrons are flowing and how the current and voltage will differ at different points on the circuits
Answer #1

10V trim 200 μF 3 Ω ΛW ΛΛΛΛΛΛΛΑ 4 Ω 4 Ω

Considering the above circuit .It has a voltage of !0 volts and the current passes through the resistor of resistance 2 ohms.according to Ohm's law V = I * R ; we can calculate the current passing through it to be I = V / R = 10 / 2 = 5 Amperes. Then the current moves on and reaches the junction where there is a parallel connection. Then current splits down into 2 paths A parallel circuit has two or more paths for current to flow through. Voltage is the same across each component of the parallel circuit. The sum of the currents through each path is equal to the total current that flows from the source.So the voltage across both the paths remain constant and the current splits according to the resistance provided.The path 1 provides and initial resistance of 3 ohm and the other a resistance of 4 ohm . So the current splits in the ratio 3 : 4. Then in the first path after passing through the 3 ohm there is a 200 μF.capacitor The capacitor stores current in it as capacitance . When a capacitor is fully charged then the potential difference across the capacitor equals the voltage source. So there will be no current in the circuit.Meanwhile the current passing through the second $ ogm resistor passes and completes the circuit .